Find least no in which when divide by 5,7,9 then it leaves remainder 1,2,3 respectively. What is logic for solving this problem.
Answer by venugopalramana(3286)

Find least no in which when divide by 5,7,9 then it leaves remainder 1,2,3 respectively. What is logic for solving this problem.
I AM NOT SURE HOW FAR YOU STUDIED ON THEORY OF NUMBERS..SO LET ME GIVE YOU A
COMMONSENSE APPOACH.......AND DOWN BELOW BASED ON THEORY OF NUMBERS
LET THE NUMBER BE .....N
N-1|5............N= 5M+1...WHERE M IS AN INTEGER.............1
N-2|7............5M+1-2|7.....5M-1|7...........5M-1=7P......2
N-3|9............5M-2|9......5M-2=9Q........................3
NOW USE EQN.3 BY GIVING Q INCREASING VALUES AND TESTING EQN 2.NOTE THAT P,Q,M,N SHOULD ALL BE INTEGERS.
I.TRY..Q=2.........5M-2=9*2=18.....M=4....BUT THIS CANT SATISFY EQN2
NEXT JUMP WILL BE TO Q=2+5=7
II.TRY.Q=7.........................M=13....NO(NOTE THAT M INCREASES IN STEPS OF 9...FOR Q INCREASIN IN STEPS OF 5)
III..Q=12..........................M=22.......NO
IV...Q=17..........................M=31......YES..5*31-1=154|7=22
SO........... N=5*31+1=156..
-------------------------------------------------------------------------
BASED ON CONGRUENCE MODULO
(5,7)=(5,9)=(7,9)=1 HENCE THERE IS A SOLUTION.
N = 1(MOD 5)=2(MOD 7) = 3(MOD 9)
N= 1+5M.....
1+5M=2(MOD 7)
5M = 1(MOD 7)....ADD 0 =14(MOD 7) TO BOTH SIDES
5M = 15 (MOD 7)
M = 3 (MOD 7)
M = 7P+3
N = 1+5(7P+3)=1+35P+15=35P+16
35P+16 = 3 (MOD 9)
35P = -13 (MOD 9).......ADD 0=153(MOD 9)
35P=140(MOD 9)
P=4(MOD 9)=9Q+4
N= 35P+16 = 35(9Q+4)+16=315Q+156
HENCE WE GET DIFFERENT SOLUTIONS FOR THE CONGRUENCIES BY PUTTING Q=0,1,2...ETC
N= 156,571,786...ETC...
LEAST VALUEOF N IS 156 AS GOT BOVE.